OpenStudy (anonymous):

The wheels on Susie's car were revolving at 40 revolutions per minute. What was the velocity of the car in centimeters per second of the wheel had a radius of 10 inches?

OpenStudy (jhannybean):

\[\sf v_{cm} = r\cdot \omega\]\[\sf \omega = \frac{40~\text{rev}}{1~\text{min}}\cdot \frac{1~\text{min}}{60~\text{sec}}=\frac{40~\text{rev}}{60~\text{sec}}\]\[\sf r= 10 ~\text{in} \cdot \frac{2.54~\text{cm}}{1~\text{in}} = (10 \cdot 2.54) ~\text{cm}\]

OpenStudy (jhannybean):

Now just plug in these values into the formula

OpenStudy (jhannybean):

you should also reduce the values as well

jimthompson5910 (jim_thompson5910):

You will also need to find the circumference C and use the idea that 1 rev corresponds to C to help find the final answer